# Impact of Data Duplication on Deep Neural Network-Based Image Classifiers: Robust vs. Standard Models

### v2

@@ Abstract
Thus, conducting a thorough analysis and evaluation of dataset quality is crucial for ensuring the development of effective data preprocessing methods. Preprocessing involves cleaning raw data through various steps including removing redundancies and outliers. The recently proposed DeepSeek language model [5], which has garnered significant attention, also highlights the importance of high-quality data in building efficient models. One concern that may hinder the efficiency of a language model is the presence of duplicate entries in text datasets, as discussed in [6] that emphasize the importance of removing duplicates before training.
Similarly, duplicate images frequently appear in real-world datasets. Consequently, a similar conclusion can be drawn that duplicates may affect machine learning models for image processing, particularly in image classification tasks. However, the impact of duplication on image classification has received limited attention. Previous studies on image classification have primarily examined the effect of training set images being duplicated on the test sets [7, 8].
We hypothesize that the existence of duplicates within the training set in image classifier DNNs will also prohibit general learning and will negatively affect efficiency, accuracy, and robustness against attacks like adversarial attacks[9, 10]. However, existing literature does not examine the impact of duplicates in training data, that leaves their effects unexplored.
In this work, we address these gaps in the literature by conducting a comprehensive study on the impact of duplicates in image classification. We examine the impact of varying levels of duplication on model generalization and training efficiency, by considering both uniform and non-uniform duplication scenarios. We also evaluate the impact of repetitive data within the dataset of an adversarially trained model [11], shedding light on its influence on robustness and accuracy.
@@ 1.1 Our Contribution
The previous works studying the effect of repetitive data within datasets indicated that duplication can increase accuracy and result in better generalization while being susceptible to memorialization. In this paper, our primary goal is obtaining a complete understanding of the impact of image duplication in the training set on generalization. We show that duplication in image classifier datasets does not always yield better generalization and can negatively affect it, especially in non-uniform duplications and adversarial settings.
Our primary goal is to present a comprehensive analysis of how duplicates affect the accuracy of image classifications. We will first conduct a theoretical analysis of duplication impact in different scenarios to better understand its effect on an image classifier. Assuming data is selected from two Gaussian distributions, we will evaluate the impact of adding duplicate data. Selecting both non-uniform and uniform duplication with regard to duplicated data point labels, our analysis shows that non-uniform duplication leads to biased decision boundaries and reduces model generalization to unseen data. Our empirical results on the CIFAR-10 dataset, when randomly selected images have been duplicated, show no considerable accuracy improvement. A similar experiment on the adversarially trained model on the CIFAR-10 dataset shows an accuracy drop when adding duplicated images. These implementations are publicly available [12]. The main contributions of this paper include:
@@
An experiment-based analysis of the impact of repetitive data, focusing on data collected from two Gaussian distributions, examining the effects of duplication when the duplicated data is either uniform or non-uniform will be given in Section 4.1.
An experimental-based analysis of the impact of duplication when data is selected from the CIFAR-10 dataset in both standard models trained on CIFAR-10 and adversarially trained model on CIFAR-10 will be given in Section 4.2.
@@ 2 Related Work
In contrast, the impact of duplication in image classification has not been systematically studied. Only a limited set of studies have focused on this issue. For example, Barz and Denzler [7] highlight a considerable volume of exact duplicate, near-duplicate, and very similar images between the test and training set in CIFAR-10 and CIFAR-100 [18] datasets.
@@ 3.2 Generalization Error in Adversarial Settings
According to [23], the vulnerability of machine learning models under adversarial attacks can be attributed to the bias-variance trade-off. The relationship between generalization and adversarial vulnerability can be analyzed using the bias-variance decomposition in the context of adversarial settings, as stated in Theorem 2[23].
@@ 4.2 Generalization Error with CIFAR-10 Dataset
In this section, we present an experimental-based analysis of the impact of data duplication when training models on the CIFAR-10 dataset. Our primary focus is to systematically analyze how randomly selected repetitive data influences classification accuracy in both standard models and adversarially trained models. We introduce varying levels of duplicated data during training and evaluate model performance under each duplication setting.
@@ 4.3 Results and Observations in Standard Model
Our key findings are summarized in Figure 3, which shows the test and training accuracy for models trained with varying levels of data duplication. The experiments reveal that without duplication, the test accuracy is 70.72%. With 30% duplication (-rate = 30), the test accuracy improves to 74.12%, reflecting a modest 3% increase. This suggests that increasing the number of duplications, which adds more data for training and reduces efficiency, does not lead to a substantial improvement in accuracy. Increasing the duplication rate from 30% to 60% results in only small improvements, increasing test accuracy from 74.12% to 75.11%. However, beyond this point, additional duplication yields diminishing returns, and in some cases, even slight performance degradation. This suggests that excessive duplication may introduce redundancy without contributing significant diversity to the training process.
@@ 4.4 Results and Observations in Robust Model
Adversarially trained models exhibit a more complex response to data duplication, where repetition does not improve robustness. Unlike standard models, which benefit from uniform repetition across classes, adversarially trained models are sensitive even to uniform duplication. Such repetition undermines adversarially trained models’ accuracy on clean data.
@@
This suggests that adding more data via duplication in adversarially trained models does not lead to a general improvement in robustness or accuracy, and in fact, it may hinder the model’s ability to classify clean and adversarial data effectively.
@@ Acknowledgment
This work was supported by the Flemish Government through the Cybersecurity Research Program with grant number: VOEWICS02 and in part by the Research Council KU Leuven projects IF/C1 From Website Fingerprinting to App Fingerprinting: Inferring private user activity from encrypted network traffic. This paper was also partially supported by the AIDE project funded by the Belgian SPF BOSA under the programme “Financing of projects for the development of artificial intelligence in Belgium” with reference number 06.40.32.33.00.10.” Sumanta Sarkar was supported by UKRI through the EPSRC grants EP/R029563/1 and EP/Y028813/1.
@@ Appendix A Theoretical Model: Gaussian Data Generation and SVM Classification
The data is generated using two Gaussian distributions with mean vectors and , and a shared covariance matrix . The probability density functions for the two distributions are given by:
@@
In the experiments, a Support Vector Machine (SVM) with a Radial Basis Function (RBF) kernel is used to compute the decision boundaries. The RBF kernel is defined as:

### v3

@@ Abstract
Thus, conducting a thorough analysis and evaluation of dataset quality is crucial for ensuring the development of effective data preprocessing methods. Preprocessing involves cleaning raw data through various steps, including removing redundancies and outliers. The recently proposed DeepSeek language model [10], which has garnered significant attention, also highlights the importance of high-quality data in building efficient models. One concern that may hinder the efficiency of a language model is the presence of duplicate entries in text datasets, as discussed in [19], which emphasizes the importance of removing duplicates before training.
Similarly, duplicate images frequently appear in real-world datasets. Consequently, a similar conclusion can be drawn that duplicates may affect ML models for image processing, particularly in image classification tasks. However, the impact of duplication on image classification has received limited attention. Previous studies on image classification have primarily examined the effect of training set images being duplicated on the test sets [6, 20].
We hypothesize that the existence of duplicates within the training set in image classifier DNNs will also prohibit general learning and will negatively affect efficiency, accuracy, and robustness against attacks like adversarial attacks[4, 15]. However, existing literature does not examine the impact of duplicates in training data, which leaves their effects unexplored.
In this work, we address these gaps in the literature by conducting a comprehensive study on the impact of duplicates in image classification. We examine the impact of varying levels of duplication on model generalization and training efficiency by considering both uniform and non-uniform duplication scenarios. We also evaluate the impact of repetitive data within the dataset of an adversarially trained model [23], shedding light on its influence on robustness and accuracy.
@@ 1.1 Our Contribution
The previous works studying the effect of repetitive data within datasets indicated that duplication can increase accuracy and result in better generalization while being susceptible to memorization. In this paper, our primary goal is to obtain a complete understanding of the impact of image duplication in the training set on generalization. We show that duplication in image classifier datasets does not always yield better generalization and can negatively affect it, especially in non-uniform duplications and adversarial settings.
Our primary goal is to present a comprehensive analysis of how duplicates affect the accuracy of image classification. We will first conduct a theoretical analysis of the impact of duplication in different scenarios. Assuming data is selected from two Gaussian distributions, we will evaluate the impact of adding duplicate data. Selecting both non-uniform and uniform duplication with regard to duplicated data point labels, our analysis shows that non-uniform duplication leads to biased decision boundaries and reduces model generalization to unseen data. Our empirical results on the CIFAR-10 dataset, when randomly selected images have been duplicated, show no considerable accuracy improvement. A similar experiment on the adversarially trained model on the CIFAR-10 dataset shows an accuracy drop when adding duplicated images. The implementations are publicly available [5]. The main contributions of this paper include:
@@
An experiment-based analysis of the impact of repetitive data, focusing on data collected from two Gaussian distributions, examining the effects of duplication when the duplicated data is either uniform or non-uniform, will be given in Section 4.1.
An experimental-based analysis of the impact of duplication when data is selected from the CIFAR-10 dataset in both standard models trained on CIFAR-10 and an adversarially trained model on CIFAR-10 will be given in Section 4.2.
@@ 2 Related Work
In contrast, the impact of duplication in image classification has not been systematically studied. Only a limited set of studies have focused on this issue. For example, Barz and Denzler [6] highlight a considerable volume of exact duplicates and near-duplicates between the test and training sets in CIFAR-10 and CIFAR-100 [16] datasets.
@@ 3.2 Generalization Error in Adversarial Settings
According to [3], the vulnerability of ML models under adversarial attacks can be attributed to the bias-variance trade-off. The relationship between generalization and adversarial vulnerability can be analyzed using the bias-variance decomposition in the context of adversarial settings, as stated in Theorem 2[3].
@@ 4.2 Generalization Error with CIFAR-10 Dataset
In this section, we present an experimental-based analysis of the impact of data duplication when training models on the CIFAR-10 dataset. Our primary focus is to systematically analyze how randomly selected repetitive data influences classification accuracy in both standard models and an adversarially trained model, which is trained under Projected Gradient Descent (PGD) attacks constrained by an norm with . We introduce varying levels of duplicated data during training and evaluate model performance under each duplication setting.
@@ 4.3 Results and Observations in Standard Model
Our key findings are summarized in Figure 3, which shows the test and training accuracy for models trained with varying levels of data duplication. The experiments reveal that without duplication, the test accuracy is 70.72%. With 30% duplication (-rate = 30), the test accuracy improves to 74.12%, reflecting a modest 3% increase. This suggests that increasing the number of duplications, which adds more data for training and reduces efficiency, does not lead to a substantial improvement in accuracy. Increasing the duplication rate from 30% to 60% results in only small improvements, increasing test accuracy from 74.12% to 75.11%. However, beyond this point, additional duplication yields diminishing returns, and in some cases, even slight performance degradation. This shows that excessive duplication may introduce redundancy without contributing significant diversity to the training process.
@@ 4.4 Results and Observations in Robust Model
Adversarially trained models exhibit a more complex response to data duplication, where repetition does not improve robustness. Unlike standard models, which benefit from uniform repetition across classes, adversarially trained models are sensitive even to uniform duplication. Such repetition undermines adversarially trained models’ accuracy.
@@
This suggests that adding duplications in adversarially trained models does not lead to a general improvement in robustness or accuracy, and in fact, it may hinder the model’s ability to classify clean and adversarial data correctly.
@@ Acknowledgment
This work was supported by the Flemish Government through the Cybersecurity Research Program with grant number: VOEWICS02 and in part by the Research Council KU Leuven projects IF/C1 From Website Fingerprinting to App Fingerprinting: Inferring private user activity from encrypted network traffic, and the AIDE project funded by the Belgian SPF BOSA under the programme “Financing of projects for the development of artificial intelligence in Belgium” with reference number 06.40.32.33.00.10.” Sumanta Sarkar was supported by UKRI through the EPSRC grants EP/T014784/1 and EP/X036669/1.
@@ Appendix A Theoretical Model: Gaussian Data Generation and SVM Classification
The data is generated using two Gaussian distributions with mean vectors and , and a shared covariance matrix . The probability density function for class where is given by:
@@
In the experiments, an SVM with a Radial Basis Function (RBF) kernel is used to compute the decision boundaries. The RBF kernel is defined as: